OpenStudy (anonymous):

Construct a 3X3 matrix with the following column space, row space and null space

OpenStudy (anonymous):

column space contains \[\left(\begin{matrix}1 \\ 1\\2\end{matrix}\right)\]

OpenStudy (anonymous):

row space contains [1 2 3]

OpenStudy (anonymous):

and nullspace contains \[\left(\begin{matrix}2 \\ -1\\0\end{matrix}\right) and \left(\begin{matrix}1 \\ 1\\-1\end{matrix}\right)\]